# Why urban heterogeneity limits the 15-minute city

### 12 Mar 2026 v1

@@ Abstract
This observation raises a fundamental question left open by existing studies Büttner et al. (2024); Omwamba et al. (2025): under what conditions is it possible, even in principle, to locate firms in space such that every worker can reach their workplace within a prescribed travel time $\tau_{0}$? We address this question using a minimal spatial framework combining three ingredients: (i) a population distributed over a city of area $A$, (ii) firms whose sizes follow a Zipf distribution with exponent $\gamma$, and (iii) an optimization problem in which firm locations are chosen to minimize the maximum commuting distance. This framework yields strict theoretical bounds on proximity, which we confront with numerical optimization and a case study of Paris and its near suburbs.
@@
To isolate the role of the journey-to-work commute, we introduce a minimal model that focuses on this single ingredient and reveals the existence of a structural lower bound on commuting time. We thus consider an idealized city of characteristic radius $R$ (typically of order $\sqrt{A/\pi}$, where $A$ is the urban area), and total employment $E$. Residents are assumed to be uniformly distributed in space unless stated otherwise. Employment is provided by $N$ firms whose sizes are heterogeneous and ranked by decreasing employment $m_{1}>m_{2}>\dots>m_{N}$. Firm sizes are assumed to follow a Zipf distribution, such that the firm of rank $r$ employs $m_{r}\propto r^{-\gamma}$ workers, where the exponent $\gamma$ controls the degree of economic inequality: $\gamma=0$ corresponds to homogeneous firms, while larger values of $\gamma$ indicate increasingly concentrated employment in a small number of large firms.
@@
where $H_{N}(\gamma)=\sum_{r=1}^{N}r^{-\gamma}$ is the $n$-th generalized harmonic number. For large $N$, $H_{N}(\gamma)\sim N^{1-\gamma}/(1-\gamma)$ for $\gamma<1$, $H_{N}(1)\sim\log N$ in the marginal case, and $H_{N}(\gamma)\to\zeta(\gamma)$ for $\gamma>1$, where $\zeta(\gamma)$ denotes the Riemann zeta function. This result Eq. 1 shows that the firm-size distribution alone determines how many workers must be served by the largest employer, independently of spatial considerations.
@@ Results
The factor $\sqrt{A/\pi}/v$ corresponds to the time required to traverse the city radius, while the effect of firm-size heterogeneity enters through the denominator $\sqrt{H_{N}(\gamma)}$. The resulting constraint depends qualitatively on the value of $\gamma$. For $\gamma>1$, the harmonic number converges to $H_{N}(\gamma)\to\zeta(\gamma)$ in the large-$N$ limit, yielding a finite asymptotic lower bound $L_{0}^{\ast}=\sqrt{A/\pi}/\sqrt{\zeta(\gamma)}$. In this regime, increasing the number of firms cannot reduce commuting distances below a fixed fraction of the city size, revealing a structural obstruction to universal short commutes. By contrast, in the homogeneous limit $\gamma=0$, one has $H_{N}(0)=N$ and the bound reduces to the purely geometric scaling $L_{0}^{\ast}=\sqrt{A/\pi}/\sqrt{N}$, which can be made arbitrarily small by increasing the number of firms.
@@
We show in Fig. 1 the result of the optimization for a representative example with $E=1000$ employees and $N=60$ firms, with firm sizes distributed according to a Zipf law with exponent $\gamma=1.5$. Starting from a random spatial configuration of firms and a random allocation of workers to firms, the simulated annealing algorithm searches for the arrangement that minimizes the maximal commuting distance $\ell_{\max}$. In the initial configuration, the maximal commute is $\ell_{\max}=1.68$, which is close to the diameter $2$ of the disk. After optimization, the maximal commuting distance decreases to $\ell_{\max}\approx 0.70$, corresponding to roughly one quarter of the system diameter. The optimized configuration exhibits a clear spatial organization: the largest firm is located near the center of the system, while smaller firms are distributed toward the periphery. This arrangement minimizes extreme commuting distances and prevents the occurrence of very long trips.
@@
Paris provides a particularly challenging case for the feasibility of a 15-minute city. Approximately $1.8$ million individuals work inside the city of Paris, but commuting flows are highly asymmetric: nearly $60\%$ (1,080,0001{,}080{,}000 individuals) of workers employed in Paris reside outside the city (data for 2017; INSEE (2022)). Among these $1.08$ million workers, about $55\%$ come from the near suburbs (the petite couronne), with the remainder residing in more distant suburbs and other regions. We therefore focus on the $1.314$ million individuals who work in Paris and live either in Paris or in its near suburbs, excluding the outer suburbs (the grande couronne), where commuting distances typically exceed $20$ km and a 15-minute constraint is unrealistic.
@@
Paris hosts approximately 406 000406\,000 establishments Greffe du Tribunal de Commerce de Paris (2018), corresponding to an average firm size of $\langle m\rangle\simeq 4.4$ employees. Rather than relying on a single estimate of $N$ for the combined area of Paris and the petite couronne (which for example would be $\sim 3\times 10^{5}$ for $4.4$ employees per firm), we vary $N$ over almost two orders of magnitude, from $10^{4}$ to $5\times 10^{5}$, encompassing any plausible value. To characterize firm-size heterogeneity, we estimate the Zipf exponent $\gamma$ using establishment-size data available at the scale of the Paris urban area. Using INSEE data for 2023 INSEE (2023), we obtain a relatively large value $\gamma\simeq 1.38$, indicating strong employment concentration. Given uncertainties in the estimation, we treat $\gamma$ as a control parameter and explore a broad range of values.
@@
The central question is whether these $N$ (with $N\in[10^{4},5\times 10^{5}]$) establishments can be spatially reorganized over Paris and the petite couronne such that all $1.314$ million workers commute less than 15 minutes. To address this question, we use the theoretical lower bound $\tau_{\min}(\gamma)$ derived above. At this scale, population density can be reasonably approximated as uniform, allowing direct application of the analytical framework.
@@
Figure 3 shows the resulting minimal commuting time $\tau_{0}(\gamma)$ as a function of firm-size heterogeneity for both walking and cycling, across a wide range of establishment counts $N$. Two main regimes emerge. For $\gamma\gtrsim 1$, the curves at different $N$ collapse onto a single line: in this regime $H_{N}(\gamma)$ is well approximated by $\zeta(\gamma)$, and the theoretical bound depends only on the area and the exponent, not on the number of firms. For $\gamma\lesssim 1$, by contrast, $H_{N}(\gamma)$ grows without bound with $N$, making $\tau_{0}$ sensitive to the precise establishment count; larger $N$ spreads workers across more, smaller catchment areas and thereby reduces the worst-case commuting distance. for Paris, the empirical firm-size distribution, however, falls in the $\gamma>1$ regime, where the bound is $N$-independent. The shaded bands quantify the additional gain from capacity flexibility $\delta$. The multiplicative reduction $\sqrt{1-\delta}$ is independent of $\gamma$ (see SI), but even for $\delta=0.3$—corresponding to a roughly $16\%$ decrease in $\tau_{0}$—the correction is far too modest to bring commuting times below the 15-minute threshold for any empirically relevant value of $\gamma$. Crucially, these results show that for a wide range of $\gamma$ values compatible with empirical estimates, including $\hat{\gamma}\simeq 1.38$, the 15-minute threshold cannot be reached—neither by increasing the number of establishments, nor by allowing substantial capacity flexibility. While increasing travel speed through cycling significantly reduces commuting times, it is generally insufficient to offset the constraint imposed by firm-size heterogeneity. These results indicate that, at the scale of Paris and its near suburbs, achieving a 15-minute city would require not only improvements in mobility but a substantial reorganisation of economic activity itself.
@@
Importantly, the framework developed here also offers a practical way to estimate the minimum achievable proximity time in a given city. When empirical data are available for the firm-size distribution (through $\gamma$, $N$, and $E$) and for the spatial extent $A$ of the urban area, the theory yields a lower bound on commuting time that no spatial rearrangement of workplaces alone can overcome. In this sense, the relevant question is not whether a city can be made into an “$x$-minute city” for a prescribed value of $x$, but rather what the smallest feasible $x$ is, given its economic structure and geometry. This perspective suggests a differentiated strategy: proximity-based planning may successfully ensure local access to daily amenities, while access to employment in heterogeneous urban economies may require complementary mobility solutions and efficient transport networks rather than purely spatial redistribution. By quantifying these structural limits, this work provides a framework for evaluating where and when proximity-based urban planning can realistically succeed.

### 16 Mar 2026 v2

@@ Abstract
This observation raises a fundamental question left open by existing studies Büttner et al. (2024); Omwamba et al. (2025): under what conditions is it possible, even in principle, to locate firms in space such that every worker can reach their workplace within a prescribed travel time $\tau_{0}$? We address this question using a minimal spatial framework combining three ingredients: (i) a population distributed over a city of area AA, (ii) firms whose sizes follow a Zipf distribution with exponent $\gamma$, and (iii) an optimization problem in which firm locations are chosen to minimize the maximum commuting distance. This framework yields strict theoretical bounds on proximity, which we confront with numerical optimization and a case study of Paris and its near suburbs.
@@
To isolate the role of the journey-to-work commute, we introduce a minimal model that focuses on this single ingredient and reveals the existence of a structural lower bound on commuting time. We thus consider an idealized city of characteristic radius $R$ (typically of order $\sqrt{A/\pi}$, where AA is the urban area), and total employment $E$. Residents are assumed to be uniformly distributed in space unless stated otherwise. Employment is provided by NN firms whose sizes are heterogeneous and ranked by decreasing employment $m_{1}>m_{2}>\dots>m_{N}$. Firm sizes are assumed to follow a Zipf distribution, such that the firm of rank $r$ employs $m_{r}\propto r^{-\gamma}$ workers, where the exponent $\gamma$ controls the degree of economic inequality: $\gamma=0$ corresponds to homogeneous firms, while larger values of $\gamma$ indicate increasingly concentrated employment in a small number of large firms.
@@
where $H_{N}(\gamma)=\sum_{r=1}^{N}r^{-\gamma}$ is the $n$-th generalized harmonic number. For large NN, $H_{N}(\gamma)\sim N^{1-\gamma}/(1-\gamma)$ for $\gamma<1$, $H_{N}(1)\sim\log N$ in the marginal case, and $H_{N}(\gamma)\to\zeta(\gamma)$ for $\gamma>1$, where $\zeta(\gamma)$ denotes the Riemann zeta function. This result Eq. 1 shows that the firm-size distribution alone determines how many workers must be served by the largest employer, independently of spatial considerations.
@@ Results
The factor $\sqrt{A/\pi}/v$ corresponds to the time required to traverse the city radius, while the effect of firm-size heterogeneity enters through the denominator $\sqrt{H_{N}(\gamma)}$. The resulting constraint depends qualitatively on the value of $\gamma$. For $\gamma>1$, the harmonic number converges to $H_{N}(\gamma)\to\zeta(\gamma)$ in the large-NN limit, yielding a finite asymptotic lower bound $L_{0}^{\ast}=\sqrt{A/\pi}/\sqrt{\zeta(\gamma)}$. In this regime, increasing the number of firms cannot reduce commuting distances below a fixed fraction of the city size, revealing a structural obstruction to universal short commutes. By contrast, in the homogeneous limit $\gamma=0$, one has $H_{N}(0)=N$ and the bound reduces to the purely geometric scaling $L_{0}^{\ast}=\sqrt{A/\pi}/\sqrt{N}$, which can be made arbitrarily small by increasing the number of firms.
@@
We show in Fig. 1 the result of the optimization for a representative example with $E=1000$ employees and $N=60$ firms, with firm sizes distributed according to a Zipf law with exponent $\gamma=1.5$. Starting from a random spatial configuration of firms and a random allocation of workers to firms, the simulated annealing algorithm searches for the arrangement that minimizes the maximal commuting distance $\ell_{\max}$. In the initial configuration, the maximal commute is $\ell_{\max}=1.68$, which is close to the diameter 22 of the disk. After optimization, the maximal commuting distance decreases to $\ell_{\max}\approx 0.70$, corresponding to roughly one quarter of the system diameter. The optimized configuration exhibits a clear spatial organization: the largest firm is located near the center of the system, while smaller firms are distributed toward the periphery. This arrangement minimizes extreme commuting distances and prevents the occurrence of very long trips.
@@
Paris provides a particularly challenging case for the feasibility of a 15-minute city. Approximately $1.8$ million individuals work inside the city of Paris, but commuting flows are highly asymmetric: nearly $60\%$ (1,080,0001{,}080{,}000 individuals) of workers employed in Paris reside outside the city (data for 2017; INSEE (2022)). Among these $1.08$ million workers, about $55\%$ come from the near suburbs (the petite couronne), with the remainder residing in more distant suburbs and other regions. We therefore focus on the $1.314$ million individuals who work in Paris and live either in Paris or in its near suburbs, excluding the outer suburbs (the grande couronne), where commuting distances typically exceed 2020 km and a 15-minute constraint is unrealistic.
@@
Paris hosts approximately 406 000406\,000 establishments Greffe du Tribunal de Commerce de Paris (2018), corresponding to an average firm size of $\langle m\rangle\simeq 4.4$ employees. Rather than relying on a single estimate of NN for the combined area of Paris and the petite couronne (which for example would be $\sim 3\times 10^{5}$ for $4.4$ employees per firm), we vary NN over almost two orders of magnitude, from $10^{4}$ to $5\times 10^{5}$, encompassing any plausible value. To characterize firm-size heterogeneity, we estimate the Zipf exponent $\gamma$ using establishment-size data available at the scale of the Paris urban area. Using INSEE data for 2023 INSEE (2023), we obtain a relatively large value $\gamma\simeq 1.38$, indicating strong employment concentration. Given uncertainties in the estimation, we treat $\gamma$ as a control parameter and explore a broad range of values.
@@
The central question is whether these NN (with $N\in[10^{4},5\times 10^{5}]$) establishments can be spatially reorganized over Paris and the petite couronne such that all $1.314$ million workers commute less than 15 minutes. To address this question, we use the theoretical lower bound $\tau_{\min}(\gamma)$ derived above. At this scale, population density can be reasonably approximated as uniform, allowing direct application of the analytical framework.
@@
Figure 3 shows the resulting minimal commuting time $\tau_{0}(\gamma)$ as a function of firm-size heterogeneity for both walking and cycling, across a wide range of establishment counts NN. Two main regimes emerge. For $\gamma\gtrsim 1$, the curves at different NN collapse onto a single line: in this regime $H_{N}(\gamma)$ is well approximated by $\zeta(\gamma)$, and the theoretical bound depends only on the area and the exponent, not on the number of firms. For $\gamma\lesssim 1$, by contrast, $H_{N}(\gamma)$ grows without bound with NN, making $\tau_{0}$ sensitive to the precise establishment count; larger NN spreads workers across more, smaller catchment areas and thereby reduces the worst-case commuting distance. for Paris, the empirical firm-size distribution, however, falls in the $\gamma>1$ regime, where the bound is NN-independent. The shaded bands quantify the additional gain from capacity flexibility $\delta$. The multiplicative reduction $\sqrt{1-\delta}$ is independent of $\gamma$ (see SI), but even for $\delta=0.3$—corresponding to a roughly $16\%$ decrease in $\tau_{0}$—the correction is far too modest to bring commuting times below the 15-minute threshold for any empirically relevant value of $\gamma$. Crucially, these results show that for a wide range of $\gamma$ values compatible with empirical estimates, including $\hat{\gamma}\simeq 1.38$, the 15-minute threshold cannot be reached—neither by increasing the number of establishments, nor by allowing substantial capacity flexibility. While increasing travel speed through cycling significantly reduces commuting times, it is generally insufficient to offset the constraint imposed by firm-size heterogeneity. These results indicate that, at the scale of Paris and its near suburbs, achieving a 15-minute city would require not only improvements in mobility but a substantial reorganisation of economic activity itself.
@@
Importantly, the framework developed here also offers a practical way to estimate the minimum achievable proximity time in a given city. When empirical data are available for the firm-size distribution (through $\gamma$, NN, and $E$) and for the spatial extent AA of the urban area, the theory yields a lower bound on commuting time that no spatial rearrangement of workplaces alone can overcome. In this sense, the relevant question is not whether a city can be made into an “$x$-minute city” for a prescribed value of $x$, but rather what the smallest feasible $x$ is, given its economic structure and geometry. This perspective suggests a differentiated strategy: proximity-based planning may successfully ensure local access to daily amenities, while access to employment in heterogeneous urban economies may require complementary mobility solutions and efficient transport networks rather than purely spatial redistribution. By quantifying these structural limits, this work provides a framework for evaluating where and when proximity-based urban planning can realistically succeed.